# Continual Distillation Learning: An Empirical Study of Knowledge Distillation in Prompt-based Continual Learning

### v2

@@ Abstract
Knowledge Distillation (KD) focuses on using a teacher model to improve a student model. Traditionally, KD is studied in an offline fashion, where a training dataset is available before learning. In this work, we introduce the problem of Continual Distillation Learning (CDL) that considers KD in the Continual Learning (CL) setup. A teacher model and a student model need to learn a sequence of tasks, and the knowledge of the teacher model will be distilled to the student to improve the student model in an online fashion. The CDL problem is valuable to study since for prompt-based continual learning methods, using a larger vision transformer (ViT) leads to better performance in continual learning. Distilling the knowledge from a large ViT to a small ViT can improve inference efficiency for prompt-based CL models. To this end, we conducted experiments to study the CDL problem with three prompt-based CL models, i.e., L2P, DualPrompt and CODA-Prompt, where we utilized logit distillation, feature distillation and prompt distillation for knowledge distillation from a teacher model to a student model. Our findings of this study can serve as baselines for future CDL work. [^1] Project website https://irvlutd.github.io/CDL/
@@ 1 Introduction
Continual Learning (CL) [30] designs models that can continuously learn new tasks without forgetting previously learned tasks. In class-incremental continual learning, data of new classes arrive sequentially, and the model needs to learn to recognize these classes sequentially during training. In testing, the model will be tested on all the seen classes. Therefore, a good CL model should learn new classes without forgetting.
Traditionally, the class-incremental setting is tackled using a rehearsal buffer [24, 17, 21, 3], which stores training examples from previous classes. During training, data from new tasks and data from the rehearsal buffer are used to update the model weights. With the help of the rehearsal buffer, the model can continually refresh its memory of previous classes. However, due to the size limit of the rehearsal buffer, these models still suffer from the catastrophic forgetting problem [22].
Recently, a new paradigm for CL based on learning for prompt receives more attentions [34, 33, 28, 31]. Instead of using rehearsal buffers, these methods focus on learning prompts for pre-trained large models such as vision transformers for continual learning. For example, L2P [34] learns a prompt pool for a set of tasks using pre-trained ViTs [6], where different tasks learn to use different prompts in the prompt pool. DualPrompt [33] learns two types of prompts, i.e., a global prompt for all tasks and task-specific prompts. By leveraging the pre-trained ViT models, these models achieve better performance compared to traditional rehearsal buffer-based methods.
We noticed that for prompt-based CL methods using ViTs, larger ViTs [6] achieve better performance. For example, ViT large is better than ViT base, and ViT base is better than ViT small, etc. Therefore, it is appealing to use large models in prompt-based CL. However, larger models introduce more computation during inference. To address this problem, we study the problem of Knowledge Distillation (KD) from large to small models in the continual learning setting. We name this problem Continual Distillation Learning (CDL).
@@
Figure 1: Comparison between three different knowledge distillation scenarios. (a) Traditional KD with an offline dataset. (b) KD in rehearsal-based CL models. (c) The CDL problem introduced in this work.
@@
The CDL problem is different from the previous knowledge distillation setup. We illustrate these differences in Fig. 1. First, in the traditional KD setup, an offline dataset is used for knowledge distillation (Fig. 1(a)). A teacher model is first trained using the dataset, and then its knowledge is distilled to a student model using the same dataset. The setup is not the continual learning setup. Second, KD is also used in the rehearsal-based CL methods such as iCaRL [24] (Fig. 1(b)). These methods maintain a rehearsal buffer of previous data. When updating the model parameters for the current task, KD ensures that the prediction on the previous data should be similar to the prediction of the old model. In this case, the old model before updating is the teacher model. Lastly, the CDL problem we study in this work is illustrated in Fig. 1(c). Given a new task in continual learning, a teacher model is first updated. Then a student model is updated based on the data of the new task and the teacher model. Consequently, a better student model can be trained by knowledge distillation from the teacher model.
To study the CDL problem, we have designed a prompt-based continual learning method that consists of a teacher model and a student model. In our method, we jointly learn the teacher model and the student model. Both models are prompt-based CL models, where the backbones are frozen and the prompts are learnable. When training data for new tasks come, the teacher model is first updated as other prompt-based CL models in the literature, and then the student model is updated.
Our method is flexible to be used with different prompt-based CL models and different mechanisms for knowledge distillation. Therefore, we conducted an empirical study of the CDL problem using three prompt-based CL models, i.e., L2P [34], DualPrompt [33] and CODA-Prompt [28]. For each model, we considered different teacher-student configurations. For example, using ViT large as the teacher model and ViT base as the student model. In addition, we studied the effect of three types of knowledge distillation mechanism in the CDL setup, i.e. logit distillation, feature distillation, and prompt distillation. We found that prompt distillation works the best, due to the prompt learning nature of these models.
@@
  • We introduce the problem of continual distillation learning (CDL) that studies knowledge distillation of a teacher model to a student model in continual learning.
  • We designed a method for CDL that can be used with different prompt-based CL models and different knowledge distillation mechanism.
  • We conducted extensive experiments and analyses of three prompt-based CL models with three types of KD mechanism on two commonly-used benchmarks for continual learning.
@@ 2 Related Work
Continual Learning. The main purpose of continual learning is to build an intelligent system to solve the problem of catastrophic forgetting  [22] in the case of incremental tasks. Different types of continual learning methods have been proposed in the literature [32]. For example, [24, 17] utilize function regularization to help the loss function. There are also weight regularization methods [15, 38, 20, 25, 27] that selectively constrain changes in network parameters and impose penalties on changes in each parameter based on its contribution. Architecture-based approaches [37, 13, 35, 7, 23] mainly focus on constructing a special model for continual learning. Replay-based approaches [24, 17, 12, 21, 1, 3, 2] use the memory buffer to store and replay old task data in learning the current task. Some recent works aim to address the issue of catastrophic forgetting without relying on rehearsal memory, which are referred as rehearsal-free methods [9, 8, 5]. The aforementioned methods do not leverage pre-trained networks. They usually learn the tasks from scratch.
Prompt-based Continual Learning. Inspired by the use of prompts in natural language processing [19], prompt-based continual learning methods employ large Vision Transformer (ViT) models [6]. These methods freeze the pre-trained backbone and shift learning to the prompts of the ViTs. For example, L2P [34] learns a prompt pool of key-prompt pairs, then selects the optimal prompt in the pool for a given input by matching the input with the keys in the pool to select the corresponding prompts. Based on this idea, DualPrompt [33] introduces the concept of using a general prompt and an expert prompt. CODA-Prompt [28] abandons the key-prompt pair selection idea. It uses a weighted sum of the prompt components to obtain the final prompt. In prompt-based methods, the size of the backbone matters. Larger backbone models achieve better performance. This motivates us to study the knowledge distillation problem for prompt-based continual learning methods, where we can distill knowledge from large models to small models in order to speed up inference time.
Knowledge Distillation (KD). The purpose of knowledge distillation is to leverage large models to assist smaller models in fully exploiting their potential. The larger model, termed the teacher, aids in training the smaller model, referred to as the student. According to the distillation location, traditional methods can be divided into logit distillation [11, 39] and feature distillation [18, 26, 4]. With the recent rapid advancements in large models, new research has directed the focus of knowledge distillation towards vision transformers [29, 36]. However, KD for prompt-based continual learning models has not yet been explored, which is the focus of this work.
@@ 3 Prerequisites
3.1 L2P
L2P [34] and DualPrompt [33] use a pre-trained ViT model as the backbone and keep the backbone frozen. Both methods use a prompt pool to encode information about tasks: $\mathbf{P}=\{P_{1},P_{2},\ldots,P_{M}\}$, where $P_{i}\in\mathbb{R}^{L_{p}\times D}$ with token length $L_{p}$ and embedding size $D$, and $M$ is the total number of prompts. In order to select prompts for different tasks, each prompt is associated with a learnable key $\{(\mathbf{k}_{1},P_{1}),(\mathbf{k}_{2},P_{2}),\ldots,(\mathbf{k}_{M},P_{M})\}$, where $\mathbf{k}_{i}\in\mathbb{R}^{D_{k}}$ and $\mathbf{K}=\{\mathbf{k}_{i}\}_{i=1}^{M}$ denotes all the keys.
@@ 3.1 L2P
Given an input image $\mathbf{x}$, a query function $q(\cdot)$ is used to encode the input image. The query $\mathbf{x}_{e}=q(\mathbf{x})\in\mathbb{R}^{D_{k}}$ then matches the key $\mathbf{k}_{i}$ with cosine similarity: $\gamma\left(q(x),\mathbf{k}_{i}\right)$, where $\gamma(\cdot,\cdot)$ denotes the cosine similarity function. Then the top $k$ prompts from the prompt pool $\mathbf{P}$ based on the establishment of key-value pairs are selected. These prompts are used in the ViT model to classify the input. In this way, given an input-label pair $(\mathbf{x},y)$, L2P optimizes the trainable prompt keys and values, and the final classification layer. The training loss function is defined as
$\min_{\mathbf{P},\mathbf{K},\phi}\mathcal{L}(g_{\phi}(f_{b}(\mathbf{x})),y)+\lambda\sum_{\mathbf{k}_{i}\in\mathbf{K}_{\mathbf{x}}}\gamma(q(\mathbf{x}),\mathbf{k}_{i}),$ (1)
@@
where $g_{\phi}$ is the classifier, $f_{b}$ is the pre-trained ViT backbone, $\mathbf{K}_{\mathbf{x}}$ denotes the selected keys for input $\mathbf{x}$, and $\mathcal{L}$ denotes the softmax cross-entropy loss for classification. The final class token of the ViT is used for the classifier. Therefore, the first part in Eq. (1) is the softmax cross-entropy loss to optimize the learnable prompt set $\mathbf{P}$, key set $\mathbf{K}$ and the classifier parameter $\phi$. The second term minimizes the distances between the selected keys and the corresponding query features $q(\mathbf{x})$.
3.2 DualPrompt
@@
The DualPrompt method [33] supplements insertable prompts based on the L2P method. The prompts are divided into G-Prompt (General) and E-Prompt (Expert). The shared G-Prompt among all tasks and the corresponding E-Prompt are attached to multiple multi-head self-attention (MSA) layers of the pre-trained transformer. The prompts in both the E-Prompt and L2P are identical, primarily serving to distinguish between different tasks. In contrast, the G-Prompt serves as a shared prompt across different tasks, mainly representing the shared information between tasks. In training, DualPrompt optimizes the G-prompt and E-prompt jointly. Please see [33] for details.
3.3 CODA-Prompt
@@
Based on DualPrompt, CODA-Prompt [28] desires an approach that is end-to-end differentiable. Without selecting the top-K key-value pairs in the prompt pool, CODA-Prompt builds the prompt components: $\mathbf{p}=\sum_{m}\alpha_{m}P_{m}$, where $\mathbf{P}=\{P_{1},P_{2},\ldots,P_{M}\}\subseteq\mathbb{R}^{L_{p}\times D\times M}$ is the set of prompt components, $M$ is the length of the set. $\alpha_{m}$ is the weight for the component $P_{m}$, which is computed using the cosine similarity of the Query-Key. Unlike L2P and Dual-Prompt, when calculating the similarity, CODA-Prompt creates a feature-selection attention scheme $\mathbf{A}$ to process the query features: $\alpha=\gamma\left(q(\mathbf{x})\odot\mathbf{A},\mathbf{K}\right)$, where $\mathbf{A}\subseteq\mathbb{R}^{D_{k}\times M}$ are learnable parameters. Finally, CODA-Prompt optimizes the prompt components $\mathbf{P}$, keys $\mathbf{K}$, attention $\mathbf{A}$ and the classifier. Additionally, orthogonality constraints are added to reduce interference between existing and new knowledge. The final total loss function is:
$\min_{\mathbf{P},\mathbf{K},\mathbf{A},\phi}\mathcal{L}(g_{\phi}(f_{b}(\mathbf{x})),y)\\ +\lambda\left(\mathcal{L}_{\text{or}}\left(\mathbf{P}\right)+\mathcal{L}_{\text{or}}\left(\mathbf{K}\right)+\mathcal{L}_{\text{or}}\left(\mathbf{A}\right)\right),$ (2)
@@
where $\mathbf{P}$, $\mathbf{K}$, $\mathbf{A}$ refer to the prompt components and corresponding keys/attention vectors during tasks, respectively. The orthogonality penalty loss is $\mathcal{L}_{\text{or}}(B)=\|BB^{T}-I\|_{2}$, where $B$ represents any matrix.
3.4 Prompt Fine-tuning
Based on the characteristics of large models, they possess outstanding feature extraction capabilities. However, due to the large model size, conventional fine-tuning approaches are not particularly effective. Current prompt-based continual learning methods employ a pre-trained Vision Transformer (ViT) as the backbone, with the backbone component frozen, which effectively preserves the internal hidden features of tasks. These methods adopt a prompt fine-tuning approach, shifting the training focus to the prompt pool. This approach only requires tuning a small portion of the model (the prompt embeddings) during the tasks, thereby effectively mitigating the issue of catastrophic forgetting. Based on the manner in which prompts are incorporated, prompt fine-tuning can be categorized into two types: prompt tuning and prefix tuning.
Figure 2: Our teacher-student model for continual distillation learning. The image is processed by the query function to generate a query, which is compared with $\mathbf{K}$ in the prompt pool. Based on different prompt extractor methods, the output prompts are divided into $\mathbf{P_{K}}$ and $\mathbf{P_{V}}$, which are inserted into the MSA layers across multiple blocks. The dashed boxes represent different types of optional distillation methods. Feature distillation uses the internal tokens output by intermediate blocks, and the class-token is used in the classification layer for logit distillation. Prompt Distillation adds an additional KD-Token and KD-Classifier to establish knowledge distillation with the teacher’s classification layer. In this case, the student model’s prediction is influenced by both the KD-Classifier and the Student-Classifier.
Prompt Tuning prepends the prompt to the input tokens. Equivalent to concatenating the prompt parameters to the input query ($\mathbf{h}_{Q}$), key ($\mathbf{h}_{K}$), and values ($\mathbf{h}_{V}$) of the MSA layer, this approach will increase the output length, where the output dimension is $\mathbb{R}^{(L+L_{p})\times D}$:
$f_{\text{Pro-T}}(P,\mathbf{h})=\text{MSA}([P;\mathbf{h}_{Q}],[P;\mathbf{h}_{K}],[P;\mathbf{h}_{V}]).$ (3)

### v3

# Continual Distillation Learning: Knowledge Distillation in Prompt-based Continual Learning

@@ Abstract
We introduce the problem of continual distillation learning (CDL) in order to use knowledge distillation (KD) to improve prompt-based continual learning (CL) models. The CDL problem is valuable to study since the use of a larger vision transformer (ViT) leads to better performance in prompt-based continual learning. The distillation of knowledge from a large ViT to a small ViT can improve the inference efficiency for prompt-based CL models. We empirically found that existing KD methods such as logit distillation and feature distillation cannot effectively improve the student model in the CDL setup. To this end, we introduce a novel method named Knowledge Distillation based on Prompts (KDP), in which globally accessible prompts specifically designed for knowledge distillation are inserted into the frozen ViT backbone of the student model. We demonstrate that our KDP method effectively enhances the distillation performance in comparison to existing KD methods in the CDL setup. [^1] Project website https://irvlutd.github.io/CDL/
@@ 1 Introduction
Continual Learning (CL) [32] designs models that can continuously learn new tasks without forgetting previously learned tasks. For example, in class-incremental continual learning [26], the data of new classes arrive sequentially, and the model needs to learn to recognize these classes sequentially during training. In testing, the model will be tested on all the seen classes. Therefore, a good CL model should learn new classes without forgetting.
With recent advances in vision models, vision transformers (ViTs) [6] have demonstrated their advantages over convolutional neural networks (CNNs). Similarly, continual learning research has evolved from traditional CNN-based methods (e.g., using ResNet [11]) to the latest prompt-based CL methods that leverage ViTs as the backbone, which now represent the state-of-the-art CL models. Unlike traditional CL models where the backbone is trained, prompt-based CL methods such as L2P [35], DualPrompt [34] and CODA-Prompt [30], utilize a pre-trained ViT as the backbone while keeping it frozen during training. Instead, they optimize a prompt pool designed for continual learning. Using pre-trained ViT models and shifting learning to prompts, these methods achieve better performance compared to CNN-based continual learning approaches such as iCaRL [26] and LWF [19].
@@
Figure 1: Comparison between two knowledge distillation scenarios. (a) Traditional KD with an offline dataset. (b) The CDL problem introduced in this work.
We noticed that for prompt-based CL methods using ViTs, larger ViTs [6] achieve better performance. For example, ViT large is better than ViT base, and ViT base is better than ViT small, etc. Therefore, it is appealing to use large models in prompt-based CL. However, larger models introduce more computation during inference. To address this limitation, we propose a new research problem: Knowledge Distillation (KD) from large to small models in prompt-based continual learning. We name this problem Continual Distillation Learning (CDL), which aims to improve the performance of small models by using large models as teachers in continual learning.
We particularly limit the scope of this work to prompt-based CL models, since we have experimentally verified that large models may not result in better performance for the traditional CNN-based CL models such as the iCaRL [26] model or the LWF [19] model (see Sec. 7 in supplementary material). These CNN-based models update the backbones during training. Deeper and larger CNNs tend to overfit new tasks, failing to fundamentally solve the catastrophic forgetting problem in continual learning.
@@
The CDL problem is different from the previous knowledge distillation setup. We illustrate the differences in Fig. 1. In the traditional KD setup [13], an offline dataset is used for knowledge distillation (Fig. 1(a)). A teacher model is first trained using the dataset, and then its knowledge is distilled to a student model using the same dataset. The setup is not the continual learning setup. The CDL problem that we study in this work is illustrated in Fig. 1(b). Given a new task in continual learning, a teacher model is first updated. Then a student model is updated based on the data of the new task and the teacher model. Consequently, a better student model can be trained by distilling knowledge from the teacher model. In general, this is a task-incremental learning process, where both the teacher and student models can only access the data for the current task at each step.
To explore the new CDL problem, we first conducted an empirical study by applying existing knowledge distillation methods, i.e., logit distillation [13, 40], feature distillation [28, 4] and distillation token [31], to three prompt-based continual learning models, i.e., L2P [35], DualPrompt [34], and CODA-Prompt [30]. For each model, we consider different teacher-student configurations. For example, we can use ViT-Large as the teacher model and ViT-Base as the student model. Since the ViT backbone is frozen in prompt-based CL, knowledge transfer from the teacher to the student can only be based on the prompt pool, where the selection of prompts relies on a key-query mechanism. However, for knowledge distillation, this mechanism is not effective. Since each task independently reselects prompt components from the prompt pool, the knowledge embedded within them from the teacher model cannot be transferred to the next task. Our experimental results also show that the performance improvements achieved by these knowledge distillation methods are not significant. Previous KD methods are not effective in the CDL setting.
To address this limitation, we propose a novel method named Knowledge Distillation based on Prompts (KDP). KDP inserts globally accessible prompts, specifically designed for knowledge distillation, into the frozen ViT backbone of the student model. We name them KD prompts. These prompts are independent of the key-query mechanism of the prompt pool. They are not limited to specific tasks and can facilitate cross-task distillation. They serve as auxiliary information to guide the learning process of the student model. Experimental results demonstrate that this novel approach effectively addresses the CDL problem and enhances distillation performance compared to previous KD methods. Furthermore, we conducted an ablation study on the total number of KD prompts inserted into the backbone. The results show that the introduction of KD prompts into each ViT block significantly improves the alignment of the feature structures between layers of the teacher model and the student model, leading to better performance of the student model.
@@
  • We introduce the new problem of continual distillation learning (CDL) that studies knowledge distillation of a teacher model to a student model in continual learning.
  • We conducted an empirical study by applying different knowledge distillation methods to three prompt-based CL approaches and identify the limitations of these distillation methods in the CDL setting.
  • To address the unique challenges of the CDL problem, we propose a novel distillation method, KDP, which outperforms previous KD approaches.
@@ 2 Related Work
2.1 Continual Learning
The main purpose of continual learning is to build an intelligent system to solve the problem of catastrophic forgetting [24] in the case of incremental tasks. Different types of continual learning methods have been proposed in the literature [33]. For example, [26, 19] utilize function regularization to help the loss function. There are also weight regularization methods [17, 39, 22, 27, 29] that selectively constrain changes in network parameters and impose penalties on changes in each parameter based on its contribution. Architecture-based approaches [38, 15, 36, 7, 25] mainly focus on constructing a special model for continual learning. Replay-based approaches [26, 19, 14, 23, 1, 3, 2] use a memory buffer to store and replay old task data in learning the current task. Some recent works aim to address the issue of catastrophic forgetting without relying on rehearsal memory, which are referred as rehearsal-free methods [10, 8, 5]. The above methods are primarily based on CNN backbones, and a larger backbone model may not achieve better performance.
2.2 Prompt-based Continual Learning
Inspired by the use of prompts in natural language processing [21], prompt-based continual learning methods employ large vision transformer (ViT) models [6]. These methods freeze the pre-trained backbone and shift learning to prompts of the ViTs. For example, L2P [35] learns a prompt pool of key-prompt pairs, then selects the optimal prompt in the pool for a given input by matching the input with the keys in the pool. Based on this idea, DualPrompt [34] introduces the concept of using a general prompt and an expert prompt. CODA-Prompt [30] abandons the key-prompt pair selection idea. It uses a weighted sum of the prompt components to obtain the final prompt. CPrompt [9] introduces two main components during training: Classifier Consistency Learning (CCL) and Prompt Consistency Learning (PCL). In prompt-based methods, the size of the backbone matters. Larger backbone models achieve better performance. This motivates us to study the knowledge distillation problem for prompt-based continual learning methods, where we can distill knowledge from large models to small models in order to improve small models with fast inference time.
2.3 Knowledge Distillation
The purpose of knowledge distillation is to leverage larger models to assist smaller models to fully exploit their potential. The larger model, termed the teacher, aids in training the smaller model, termed the student. Traditional KD methods can be classified into logit distillation [13, 40] and feature distillation [20, 28, 4]. With recent rapid advancements in large models, new research has directed the focus of knowledge distillation towards vision transformers [31, 37]. However, KD for prompt-based continual learning models has not yet been explored, which is the focus of this work.
@@ 3 Prerequisites
3.1 Continual Learning Setting
In continual learning, a model is required to learn a sequence of tasks, where the data from these tasks arrive on time. We denote a sequence of tasks as $\mathcal{D}=\{\mathcal{D}_{1},\ldots,\mathcal{D}_{T}\}$, where $T$ is the number of tasks. The $t$th task $\mathcal{D}_{t}=\{(\mathbf{x}_{i}^{t},y_{i}^{t})\}_{i=1}^{n_{t}}$ consists of pairs of input sample $\mathbf{x}_{i}^{t}\in\mathcal{X}$ and its label $y_{i}^{t}\in\mathcal{Y}$, where $n^{t}$ is the number of samples for the $t$th task. In this work, we consider class-incremental learning, where each task consists of a fixed number of non-overlapping classes. In training, a model learns these tasks one by one. Data from the previous tasks are not available anymore when training future tasks. In testing, the model is evaluated by testing samples from all classes.
Figure 2: Our teacher-student model for continual distillation learning. The green and red dashed boxes represent different types of optional distillation methods. The blue dashed box represents our KDP method, where the components inside are the trainable modules to be added.
3.2 Prompt-based Methods
Prompt-based continual learning methods use pre-trained ViT backbones and freeze the backbones during training. Learning is shifted to trainable prompts.
@@ 3.2 Prompt-based Methods
L2P: L2P [35] uses a prompt pool to encode information about tasks: $\mathcal{P}=\{\mathbf{P}_{1},\mathbf{P}_{2},\ldots,\mathbf{P}_{M}\}$, where $\mathbf{P}_{i}\in\mathbb{R}^{L_{p}\times D}$ with token length $L_{p}$ and embedding size $D$, and $M$ is the total number of prompt components. To select prompts for different tasks, each prompt component is associated with a learnable key $\{(\mathbf{k}_{1},\mathbf{P}_{1}),(\mathbf{k}_{2},\mathbf{P}_{2}),\ldots,(\mathbf{k}_{M},\mathbf{P}_{M})\}$, where $\mathbf{k}_{i}\in\mathbb{R}^{D_{k}}$ with embedding size $D_{k}$ and $\mathcal{K}=\{\mathbf{k}_{i}\}_{i=1}^{M}$ denotes all the keys. Given an input image $\mathbf{x}\in\mathcal{X}$, a query function $q(\cdot)$ is used to encode the input image. The query $\mathbf{x}_{e}=q(\mathbf{x})\in\mathbb{R}^{D_{k}}$ then matches the key $\mathbf{k}_{i}$ with cosine similarity: $\gamma\left(q(\mathbf{x}),\mathbf{k}_{i}\right)$, where $\gamma(\cdot,\cdot)$ denotes the cosine similarity function. Then the top $K$ prompt components from the prompt pool $\mathcal{P}$ based on the establishment of key-value pairs are selected:
$\mathbf{P}_{\mathbf{x}}=\text{TopK}(\mathbf{P}_{1},\mathbf{P}_{2},\ldots,\mathbf{P}_{M}),$ (1)
where TopK indicates the selection function based on cosine similarity. $\mathbf{P}_{\mathbf{x}}\in\mathbb{R}^{K\times L_{p}\times D}$ is the final set of prompt components extracted from the prompt pool to assist continual learning. The training loss function for a pair $(\mathbf{x},y)$ is defined as:
$\min_{\mathcal{P},\mathcal{K},\phi}\mathcal{L}(g_{\phi}(f_{b}(\mathbf{x})),y)+\lambda\sum_{\mathbf{k}_{i}\in\mathbf{K}_{\mathbf{x}}}\gamma(q(\mathbf{x}),\mathbf{k}_{i}),$ (2)
@@
where $g_{\phi}$ is the classifier with parameter $\phi$. $f_{b}$ is the pre-trained ViT backbone, which includes the selected prompt components $\mathbf{P}_{\mathbf{x}}$ for continual learning. $\mathbf{K}_{\mathbf{x}}$ denotes the selected keys for input $\mathbf{x}$, and $\mathcal{L}$ denotes the softmax cross-entropy loss for classification. The final class token of the ViT is used for the classifier. Therefore, the first term in Eq. (2) is the softmax cross-entropy loss to optimize the learnable prompt set $\mathcal{P}$ and the classifier parameter $\phi$. The second term learns the key set $\mathcal{K}$ by minimizing the distances between the selected keys and the corresponding query features $q(\mathbf{x})$.
@@
DualPrompt: The DualPrompt method [34] supplements insertable prompts based on the L2P [35] method. The prompts are divided into G-Prompt (General) and E-Prompt (Expert). The shared G-Prompt among all tasks and the corresponding E-Prompt are attached to multiple multi-head self-attention (MSA) layers of the pre-trained transformer. The prompts in both the E-Prompt and L2P are identical, primarily serving to distinguish between different tasks. In contrast, the G-Prompt mainly represents the shared information between tasks. In training, DualPrompt optimizes the G-prompt and E-prompt jointly.
CODA-Prompt: Instead of selecting the top $K$ key-value pairs in the prompt pool, CODA-Prompt [30] uses a weighted summation over the prompt components to compute the learnable prompt parameter for an input $\mathbf{x}$:
$\mathbf{P}_{\mathbf{x}}=\sum_{i=1}^{M}\alpha_{i}\mathbf{P}_{i},$ (3)
@@
where $\alpha_{i}$ is the weight for the prompt component $\mathbf{P}_{i}$, which is computed using the cosine similarity of the query and key. Unlike L2P and DualPrompt, when calculating the similarity, CODA-Prompt creates a feature-selection attention scheme $\mathcal{A}=\{\mathbf{A}_{1},\ldots,\mathbf{A}_{M}\}$ for $M$ prompt components to process the query features, where $\mathbf{A}_{i}\in\mathbb{R}^{D_{k}},i=1,\ldots,M$. The weight is computed as $\alpha_{i}=\gamma\left(q(\mathbf{x})\odot\mathbf{A}_{i},\mathbf{k}_{i}\right)$, for $i=1,\ldots,M$. Finally, CODA-Prompt optimizes the prompt components $\mathcal{P}$, keys $\mathcal{K}$, attention $\mathcal{A}$ and the classifier. Additionally, orthogonality constraints are added to reduce interference between existing and new knowledge. The final loss function is:
$\min_{\mathcal{P},\mathcal{K},\mathcal{A},\phi}\mathcal{L}(g_{\phi}(f_{b}(\mathbf{x})),y)\\ +\lambda\left(\mathcal{L}_{\text{or}}\left(\mathcal{P}\right)+\mathcal{L}_{\text{or}}\left(\mathcal{K}\right)+\mathcal{L}_{\text{or}}\left(\mathcal{A}\right)\right),$ (4)
@@
where $\mathcal{P}$, $\mathcal{K}$, $\mathcal{A}$ refer to the prompt set and corresponding keys and attention vectors during tasks, respectively. The orthogonality penalty loss is defined as $\mathcal{L}_{\text{or}}(B)=\|BB^{T}-I\|_{2}$, where $B$ represents any matrix and $I$ is the identity matrix.
3.3 Prompting with Prefix-Tuning